OpenStudy (anonymous):

[4 - (3c - 1)][6 - ( 3c - 1)]

OpenStudy (anonymous):

is this correct? 9 c^2 - 36c + 35

OpenStudy (poofypenguin):

=[4 - (3c - 1)][6 - ( 3c - 1)] =[4-3c+1][6-3c+1] =[5-3c][7-3c] then distribute in with FOIL method... hold on a sec i'll give it a try and see if it matches your answer

OpenStudy (poofypenguin):

=[5-3c][7-3c] =35-15c-35c+9c^2 =35-50c+9c^2 i've gotten a -50c term instead of your -36c team... but i might have messed up... let me check again.

OpenStudy (poofypenguin):

Whoops... i made a mistake =[5-3c][7-3c] =35-15c-21c+9c^2 =35-36c+9c^2 Sorry about that. So yes, your initial answer is correct.